40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 26 01:00:23 2017, gautam, Update, General, DRMI locking  DRMI_1f_26April2017.pngIMG_7373.JPGIMG_7372.JPG
    Reply  Fri Apr 28 02:04:36 2017, gautam, Update, General, DRMI locking  20170428_DRMI.pdf
       Reply  Fri Apr 28 13:56:26 2017, rana, Update, General, DRMI locking  
          Reply  Mon May 1 16:29:51 2017, gautam, Update, General, DRMI locking  LSC_sensingNoise.pdfIMG_7378.JPG
             Reply  Mon May 1 21:45:54 2017, ericq, Update, General, DRMI locking  
                Reply  Wed May 3 18:45:45 2017, rana, Update, General, DRMI locking  
                   Reply  Thu May 4 19:03:15 2017, gautam, Update, General, DRMI locking - preliminary MICH NB DRMI_noArms_April30.pdfMICH_OLTF.pdfC1NB_disp_40m_MICH_NB_30_April_2017.pdf40m_REFL_RFPDs_efficiency.pdf
                      Reply  Fri May 5 10:13:02 2017, ericq, Update, General, MICH NB questions 
                         Reply  Fri May 5 12:10:53 2017, gautam, Update, General, MICH NB questions C1NB_disp_40m_MICH_NB_30_April_2017.pdf
                            Reply  Sat May 6 21:52:11 2017, rana, Update, General, MICH NB questions 
                               Reply  Wed May 10 01:56:06 2017, gautam, Update, General, MICH NB - OL coupling OL_BS_ITM_comp.pdfC1NB_disp_40m_MICH_NB_8_May_2017.pdfC1NB_disp_40m_MICH_NB_10_May_2017.pdf
                                  Reply  Wed May 10 16:53:38 2017, rana, Update, General, MICH NB - OL coupling 
                                  Reply  Wed May 10 17:17:05 2017, gautam, Update, General, DAC / Coil Driver noise actuation.jpg
                                     Reply  Wed May 10 17:46:44 2017, gautam, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                        Reply  Thu May 11 09:45:46 2017, rana, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                           Reply  Thu May 11 18:59:22 2017, gautam, Update, General, SRM coil driver + dewhite board initial survey D010001-B_40m.pdfD000183-C8_40m.pdf
                                              Reply  Fri May 12 01:36:04 2017, gautam, Update, General, SRM coil driver + dewhite board LISO modeling SRM_bypass_plus_CoilDriver.pdfliso.zip
                                                 Reply  Fri May 12 12:34:55 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                    Reply  Fri May 12 18:50:08 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                       Reply  Mon May 22 13:37:01 2017, gautam, Update, General, DAC noise estimate DAC_noise_model.pdfC1NB_disp_40m_MICH_NB_22_May_2017.pdf
                                                          Reply  Tue May 23 22:58:23 2017, gautam, Update, General, De-Whitening board noises ITMX_deWhite_ch1_noise.pdf
                                                             Reply  Thu May 25 19:27:29 2017, gautam, Update, General, Coil driver board noises  coilDriverNoises.pdf
                                                                Reply  Mon May 29 16:47:38 2017, gautam, Update, General, Coil driver boards reinstalled deWhitening_consolidated.pdfdeWhitening_measurements.zipETMY_OL.png
                                                                   Reply  Tue May 30 16:02:59 2017, gautam, Update, General, Coil driver boards reinstalled OL_noiseInjection.pdf
                                                                      Reply  Thu Jun 1 00:10:15 2017, gautam, Update, General, Coil driver boards reinstalled 
       Reply  Fri Apr 28 19:32:06 2017, gautam, Update, General, DRMI locking - PRCL angular FF PRC_angularFF.pdfPRC_TFs.pdf
Message ID: 13019     Entry time: Tue May 30 16:02:59 2017     In reply to: 13017     Reply to this: 13026
Author: gautam 
Type: Update 
Category: General 
Subject: Coil driver boards reinstalled 

I think the reason I am unable to engage the de-whitening is that the OL loop is injecting a ton of control noise - see Attachment #1. With the OL loop off (i.e. just local damping loops engaged for the ITMs), the RMS control signal at 100Hz is ~6 orders of magnitude (!) lower than with the OL loop on. So turning on the whitening was just railing the DAC I guess (since the whitening has something like 60dB gain at 100Hz).

The Oplev loops for the ITMs use an "Ellip15" low-pass filter to do the roll-off (2nd order Elliptic low pass filter with 15dB stopband atten and 2dB ripple). I confirmed that if I disable the OL loops, I was able to turn on the whitening for ITMY smoothly.

Now that the ETMY OL HeNe has been replaced, I restored alignment of the IFO. Both arms lock fine (I was also able to engage the ITMY Coil Driver whitening smoothly with the arm locked). However, something funny is going on with ASS - running the dither seems to inject huge offsets into the ITMY pit and yaw such that it almost immediately breaks the lock. This probably has to do with some EPICS values not being reset correctly since the recent slow-machine restarts (for instance, the c1iscaux restart caused all the LSC RFPD whitening gains to be reset to random values, I had to burt-restore the POX11 and POY11 values before I could get the arms to lock), I will have to investigate further.

GV edit 2pm 31 May: After talking to Koji at the meeting, I realized I did not specify what channel the attached spectra are for - it is  C1:SUS-ITMY_ULCOIL_OUT.

Quote:
 

But today I was not able to engage this transition smoothly, the optic seems to be getting kicked around when I engage the whitening. I will need to investigate this further. 


Unrelated to this work: the ETMY Oplev HeNe is dead (see Attachment #3). I thought we had just replaced this laser a couple of months ago - what is the expected lifetime of these? Perhaps the power supply at the Y-end is wonky and somehow damaging the HeNe heads?

 

Attachment 1: OL_noiseInjection.pdf  40 kB  | Hide | Hide all
OL_noiseInjection.pdf
ELOG V3.1.3-